# Unveiling the Nature and Fate of the Almost-Dark Cloud AGC 226178 through Mapping

### v1

@@ 4.2 The nature of AGC 226178: A stripped, disintegrating gas cloud or a mini-halo?
4.2.1 A former dark galaxy embedded in mini-halo?
Considering the potential contribution from non-baryonic matter, here we briefly discuss the possibility that AGC 226178 was a genuine dark galaxy embedded within a mini dark matter halo and its star formation was triggered in the past few hundred Myr after falling into the Virgo cluster. In the mini-halo scenario, the significant metallicity of AGC 226178 could be attained through mixing with the hot ICM (e.g., Tonnesen & Bryan 2021).
@@ 4.2.1 A former dark galaxy embedded in mini-halo?
Two populations of dark galaxies (without star formation) were explored by Benítez-Llambay et al. (2013) and Benítez-Llambay et al. (2017) with cosmological hydrodynamical simulations. Specifically, the two populations are coined as REionization-Limited Clouds (RELHICs) and COSmic WEB Stripped systems (COSWEBs). RELHICs are minihaloes that inhabit low-density environment and remain below the star formation threshold but are massive enough to retain baryons against photoevaporation by the cosmic UV background. COSWEBs are minihaloes that are in high-density environment and have been deprived of all baryons due to ram pressure stripping of the cosmic web. The halo mass range of these dark galaxies is and the baryonic mass 108 (Benítez-Llambay et al. 2017). Obviously, RELHICs set an upper limit for the halo mass and gas mass of these two populations of dark galaxies.
@@ 4.2.2 The survivability in the Virgo cluster: a test of ram pressure stripping
To test the response to ram pressure, the restoring pressure from the self gravity of the disc is compared at each location within the disc with the ram pressure strength from the cluster. If the ram pressure is stronger than the maximum restoring pressure found at a particular disc radius, then the gas at that radius can be assumed to be stripped. The restoring pressure that acts on the gas disc at each radius is calculated as the product of the gas disc surface density at a chosen cylindrical radius multiplied by the gravitational acceleration out of the plane of the disc (assuming face-on-ram pressure stripping). For the ram pressure from the cluster, we assume an orbital velocity of 500 (the velocity difference of the galaxy with respect to the systemic velocity of the Virgo cluster’s brightest central galaxy, Kashibadze et al. 2020 ), and a beta model for the Virgo ICM with =0.5, =50 kpc and (Model C1 from Roediger & Brüggen 2007)
@@
By assuming a thin disc, the restoring pressure from the self gravity of the disc is maximised. In Fig. 13, we plot the maximum restoring pressure as a function of the cylindrical disc radius. Since the projected distance of AGC 226178 from the cluster centre (1.25 Mpc) represents a lower limit to its true three-dimensional distance, we calculate the ram pressure assuming cluster-centric distances of 1.25, 1.75, and 2.25 Mpc, respectively. As can be seen, the restoring pressure from the disc is insufficient to maintain an extended distribution as observed in AGC 226178 ( 2 kpc in radius) for cluster-centric distances up to 2.25 Mpc. To maintain its current size, the cloud has be to at a cluster-centric distance far beyond 3 the virial radius of the Virgo cluster ( 1.0 Mpc). We also test a gas disc that is double the observed mass of AGC 226178 (). This allows a larger and more extended gas disc to survive the ram pressure. But to allow for a gas disc as extended as AGC 226178, the cluster-centric distance should be at least 3 the virial radius of Virgo.
@@
Our toy model presented above neglects the potentially relevant thermal (or confinement) pressure exerted by the ICM. However, several studies suggest that confinement pressure from the ICM could help stabilise such systems. Burkhart & Loeb (2016) proposed that clouds may reach pressure equilibrium with the external ICM, and hydrodynamical simulations by Bellazzini et al. (2018) and Calura et al. (2020) further demonstrated that, under confinement pressure, clouds may survive for timescales of 1 Gyr. While these simulations provide valuable insights into the stability of these intriguing systems, they remain far from fully realistic. For example, they assumed a low-to-moderate velocity of the clouds relative to the ICM, such that the ram pressure—a key environmental factor shown to generally dominate over confinement pressure in galaxy clusters (Bahé et al. 2012)—is less important than thermal pressure. For AGC 226178, if assuming a Virgo-centric distance of 1.7 Mpc (where the average density of the ICM is 10-28 g cm-3), a mean molecular weight of 0.6 for the hot ICM, an ICM temperature of 107 K, and a relative velocity of 500 km s-1, we estimate a ram pressure of 2.510-13 erg cm-3 and a thermal confinement pressure of 1.410-13 erg cm-3, suggesting a ram-to-thermal pressure ratio of 1.8. The three-dimensional relative velocity is likely larger than 500 km s-1, so the pressure ratio should be even higher. Therefore, while thermal confinement may contribute to the apparent stability of AGC 226178, it is unlikely to be the dominant factor. If the object resides within the cluster environment, its continued existence would therefore require a substantial reservoir of unaccounted baryonic mass.
@@
Recent wide-field searches for isolated blue stellar blobs toward the Virgo cluster by Dey et al. (2025) suggest that such stellar systems are relatively common and tend to concentrate in regions with significant X-ray emission from the ICM, while generally avoiding the central core. This apparent distribution was interpreted as evidence that these objects have passed peri-center and are now moving outwards along the elongated orbits. However, a more natural explanation is that many of these systems may actually lie far beyond the cluster in three-dimensional space but are projected onto the Virgo region. Indeed, phase-space analyses show that the interloper fraction increases significantly with projected radius from cluster centers, and many of the objects identified by Dey et al. (2025) fall into high-probability interloper zones—especially beyond 1–2 (Mamon et al. 2010). Therefore, while the location of AGC226178 is consistent with that of the Dey et al. (2025) sample, the uncertainty introduced by projection effects implies that its cluster membership cannot be definitively established from phase-space arguments alone. Nevertheless, if AGC226178 is indeed a Virgo member, environmental processes such as ram pressure stripping could plausibly contribute to the disturbed morphology.
@@ 5 Conclusions
We have presented a comprehensive analysis of 21-cm line properties of an almost dark galaxy candidate – AGC 226178. The data includes VLA multi-configuration mapping and FAST OTF mapping, obtained recently as part of the AVID project. Our VLA observations resolve the distribution in AGC 226178 down to sub-kpc scales, and our FAST observation achieve a 3 column density sensitivity of 4.81017 cm-2 for a 20 km s-1 linewidth.
The FAST observation of AGC 226178 reveals a short tail extending towards the lower velocity end in the south-west direction that points towards (in projection) the nearby dwarf galaxy VCC 2034, which has been considered the most probable point of origin of AGC 226178 in the recent literature. However, VCC 2034 exhibits a line-of-sight stripping signature that is inconsistent with an event that would have resulted in the formation of AGC 226178. The observed properties of the gas in AGC 226178 and its projected neighbouring galaxies rule out any physical association between them.
@@
6 Acknowledgements
We sincerely appreciate the anonymous referee’s highly favorable and constructive report on our work. We acknowledge support from the National Key Research and Development Program of China (grant No. 2023YFA1608100), and from the NSFC (grant Nos. 12122303, 11973039, and 12173079). This work is also supported by the China Manned Space Project (grant Nos. CMS-CSST-2021-B02, CSST-2021-A06 and CMS-CSST-2021-A07). We also acknowledge support from the CAS Pioneer Hundred Talents Program, the Strategic Priority Research Program of Chinese Academy of Sciences (grant No. XDB 41000000), and the Cyrus Chun Ying Tang Foundations.
@@ 6 Acknowledgements
This work made use of data from FAST (Five-hundred-meter Aperture Spherical radio Telescope)(https://cstr.cn/31116.02.FAST). FAST is a Chinese national mega-science facility, operated by National Astronomical Observatories, Chinese Academy of Sciences. We thank Jing Wang and Dong Yang (Kavli Institute for Astronomy and Astrophysics, Peking University, Beijing 100871, People’s Republic of China) for their assistance in developing the FAST observing strategy.
SHOH acknowledges a support from the National Research Foundation of Korea (NRF) grant funded by the Korea government (Ministry of Science and ICT: MSIT) (RS2022-00197685, RS-2023-00243222).
RS acknowledges financial support from FONDECYT Regular 2023 project No. 1230441 and also gratefully acknowledges financial support from ANID - MILENIO NCN2024112
Appendix A1 PV Diagram of VLA Observation for AGC 226178

### v2

@@ 4.2 The nature of AGC 226178: A stripped disintegrating gas cloud vs a mini-halo
4.2.1 A possible former dark galaxy embedded in mini-halo
Considering the potential contribution from non-baryonic matter, here we briefly discuss the possibility that AGC 226178 was a genuine dark galaxy embedded within a mini dark matter halo and its star formation was triggered in the past few hundred million years after falling into the Virgo cluster. In the mini-halo scenario, the significant metallicity of AGC 226178 could be attained through mixing with the hot ICM (e.g. Tonnesen & Bryan 2021).
@@ 4.2.1 A possible former dark galaxy embedded in mini-halo
Two populations of dark galaxies (without star formation) were explored by Ben´ıtez-Llambay et al. (2013) and Ben´ıtez-Llambay et al. (2017) with cosmological hydrodynamical simulations. Specifically, the two populations are denoted REionization-Limited Clouds (RELHICs) and COSmic WEB Stripped systems (COSWEBs). RELHICs are minihaloes that inhabit low-density environment and remain below the star formation threshold but are massive enough to retain baryons against photoevaporation by the cosmic UV background. COSWEBs are minihaloes that are in high-density environment and have been deprived of all baryons due to ram pressure stripping of the cosmic web. The halo mass range of these dark galaxies is and the baryonic mass 108 (Ben´ıtez-Llambay et al. 2017). Obviously, RELHICs set an upper limit for the halo mass and gas mass of these two populations of dark galaxies.
@@ 4.2.2 The survivability in the Virgo cluster: A test of ram pressure stripping
To test the response to ram pressure, the restoring pressure from the self-gravity of the disc is compared at each location within the disc with the ram pressure strength from the cluster. If the ram pressure is stronger than the maximum restoring pressure found at a particular disc radius, then the gas at that radius can be assumed to be stripped. The restoring pressure that acts on the gas disc at each radius is calculated as the product of the gas disc surface density at a chosen cylindrical radius multiplied by the gravitational acceleration out of the plane of the disc (assuming face-on-ram pressure stripping). For the ram pressure from the cluster, we assume an orbital velocity of 500 (the velocity difference of the galaxy with respect to the systemic velocity of the Virgo cluster’s brightest central galaxy, Kashibadze et al. 2020 ), and a beta model for the Virgo ICM with =0.5, =50 kpc and (Model C1 from Roediger & Brüggen 2007)
@@
By assuming a thin disc, the restoring pressure from the self-gravity of the disc is maximized. In Fig. 13, we plot the maximum restoring pressure as a function of the cylindrical disc radius. Since the projected distance of AGC 226178 from the cluster centre (1.25 Mpc) represents a lower limit to its true three-dimensional distance, we calculate the ram pressure assuming cluster-centric distances of 1.25, 1.75, and 2.25 Mpc, respectively. As can be seen, the restoring pressure from the disc is insufficient to maintain an extended distribution as observed in AGC 226178 ( 2 kpc in radius) for cluster-centric distances up to 2.25 Mpc. To maintain its current size, the cloud has be to at a cluster-centric distance far beyond 3 the virial radius of the Virgo cluster ( 1.0 Mpc). We also test a gas disc that is double the observed mass of AGC 226178 (). This allows a larger and more extended gas disc to survive the ram pressure. But to allow for a gas disc as extended as AGC 226178, the cluster-centric distance should be at least three times the virial radius of Virgo.
@@
Our toy model presented above neglects the potentially relevant thermal (or confinement) pressure exerted by the ICM. However, several studies suggest that confinement pressure from the ICM could help stabilize such systems. Burkhart & Loeb (2016) proposed that clouds may reach pressure equilibrium with the external ICM, and hydrodynamical simulations by Bellazzini et al. (2018) and Calura et al. (2020) further demonstrated that, under confinement pressure, clouds may survive for timescales of 1 Gyr. While these simulations provide valuable insights into the stability of these intriguing systems, they remain far from fully realistic. For example, they assumed a low to moderate velocityaa of the clouds relative to the ICM, such that the ram pressure—a key environmental factor shown to generally dominate over confinement pressure in galaxy clusters (Bahé et al. 2012)—is less important than thermal pressure. For AGC 226178, if assuming a Virgo-centric distance of 1.7 Mpc (where the average density of the ICM is 10-28 g cm-3), a mean molecular weight of 0.6 for the hot ICM, an ICM temperature of 107 K, and a relative velocity of 500 km s-1, we estimate a ram pressure of 2.510-13 erg cm-3 and a thermal confinement pressure of 1.410-13 erg cm-3, suggesting a ram-to-thermal pressure ratio of 1.8. The three-dimensional relative velocity is likely larger than 500 km s-1, so the pressure ratio should be even higher. Therefore, while thermal confinement may contribute to the apparent stability of AGC 226178, it is unlikely to be the dominant factor. If the object resides within the cluster environment, its continued existence would therefore require a substantial reservoir of unaccounted baryonic mass.
@@
Recent wide-field searches for isolated blue stellar blobs towards the Virgo cluster by Dey et al. (2025) suggest that such stellar systems are relatively common and tend to concentrate in regions with significant X-ray emission from the ICM, while generally avoiding the central core. This apparent distribution was interpreted as evidence that these objects have passed pericentre and are now moving outwards along the elongated orbits. However, a more natural explanation is that many of these systems may actually lie far beyond the cluster in three-dimensional space but are projected onto the Virgo region. Indeed, phase-space analyses show that the interloper fraction increases significantly with projected radius from cluster centres, and many of the objects identified by Dey et al. (2025) fall into high-probability interloper zones—especially beyond 1–2 (Mamon et al. 2010). Therefore, while the location of AGC226178 is consistent with that of the Dey et al. (2025) sample, the uncertainty introduced by projection effects implies that its cluster membership cannot be definitively established from phase-space arguments alone. Nevertheless, if AGC226178 is indeed a Virgo member, environmental processes such as ram pressure stripping could plausibly contribute to the disturbed morphology.
@@ 5 Conclusions
We have presented a comprehensive analysis of 21-cm line properties of an almost dark galaxy candidate – AGC 226178. The data includes VLA multi-configuration mapping and FAST OTF mapping, obtained recently as part of the AVID project. Our VLA observations resolve the distribution in AGC 226178 down to subkiloparsec scales, and our FAST observation achieve a 3 column density sensitivity of 4.81017 cm-2 for a 20 km s-1 linewidth.
The FAST observation of AGC 226178 reveals a short tail extending towards the lower-velocity end in the south-west direction that points towards (in projection) the nearby dwarf galaxy VCC 2034, which has been considered the most probable point of origin of AGC 226178 in the recent literature. However, VCC 2034 exhibits a line-of-sight stripping signature that is inconsistent with an event that would have resulted in the formation of AGC 226178. The observed properties of the gas in AGC 226178 and its projected neighbouring galaxies rule out any physical association between them.
@@
Acknowledgements.
@@
We sincerely appreciate the anonymous referee’s highly favorable and constructive report on our work. We thank Dr. Zhijie Qu for kindly pointing out an error in Figure 12 in the first arXiv version of the manuscript. We acknowledge support from the National Key Research and Development Program of China (grant No. 2023YFA1608100), and from the NSFC (grant Nos. 12122303, 11973039, and 12173079). This work is also supported by the China Manned Space Project (grant Nos. CMS-CSST-2021-B02, CSST-2021-A06 and CMS-CSST-2021-A07). We also acknowledge support from the CAS Pioneer Hundred Talents Program, the Strategic Priority Research Program of Chinese Academy of Sciences (grant No. XDB 41000000), and the Cyrus Chun Ying Tang Foundations. This work made use of data from FAST (Five-hundred-meter Aperture Spherical radio Telescope)(https://cstr.cn/31116.02.FAST). FAST is a Chinese national mega-science facility, operated by National Astronomical Observatories, Chinese Academy of Sciences. We thank Jing Wang and Dong Yang (Kavli Institute for Astronomy and Astrophysics, Peking University, Beijing 100871, People’s Republic of China) for their assistance in developing the FAST observing strategy. SHOH acknowledges a support from the National Research Foundation of Korea (NRF) grant funded by the Korea government (Ministry of Science and ICT: MSIT) (RS2022-00197685, RS-2023-00243222). RS acknowledges financial support from FONDECYT Regular 2023 project No. 1230441 and also gratefully acknowledges financial support from ANID - MILENIO NCN2024112References
- Bacon et al. (2014) Bacon, R., Vernet, J., Borisova, E., et al. 2014, The Messenger, 157, 13
- Bahé et al. (2012) Bahé, Y. M., McCarthy, I. G., Crain, R. A., & Theuns, T. 2012, MNRAS, 424, 1179
- Bellazzini et al. (2018) Bellazzini, M., Armillotta, L., Perina, S., et al. 2018, MNRAS, 476, 4565
- Bellazzini et al. (2022) Bellazzini, M., Magrini, L., Jones, M. G., et al. 2022, ApJ, 935, 50
- Ben´ıtez-Llambay et al. (2013) Benítez-Llambay, A., Navarro, J. F., Abadi, M. G., et al. 2013, ApJ, 763, L41
- Ben´ıtez-Llambay et al. (2017) Benítez-Llambay, A., Navarro, J. F., Frenk, C. S., et al. 2017, MNRAS, 465, 3913
- Bertoldi & McKee (1992) Bertoldi, F. & McKee, C. F. 1992, ApJ, 395, 140
- Bode et al. (2001) Bode, P., Ostriker, J. P., & Turok, N. 2001, ApJ, 556, 93
- Boselli et al. (2009) Boselli, A., Boissier, S., Cortese, L., et al. 2009, ApJ, 706, 1527
- Boselli et al. (2018a) Boselli, A., Fossati, M., Cuillandre, J. C., et al. 2018a, A&A, 615, A114
- Boselli et al. (2018b) Boselli, A., Fossati, M., Ferrarese, L., et al. 2018b, A&A, 614, A56
- Boselli et al. (2022) Boselli, A., Fossati, M., & Sun, M. 2022, A&A Rev., 30, 3
- Bournaud et al. (2007) Bournaud, F., Duc, P.-A., Brinks, E., et al. 2007, Science, 316, 1166
- Broeils & Rhee (1997) Broeils, A. H. & Rhee, M. H. 1997, A&A, 324, 877
- Burkhart & Loeb (2016) Burkhart, B. & Loeb, A. 2016, ApJ, 824, L7
- Calura et al. (2020) Calura, F., Bellazzini, M., & D’Ercole, A. 2020, MNRAS, 499, 5873
- Cannon et al. (2015) Cannon, J. M., Martinkus, C. P., Leisman, L., et al. 2015, AJ, 149, 72
- Cantiello et al. (2024) Cantiello, M., Blakeslee, J. P., Ferrarese, L., et al. 2024, ApJ, 966, 145
- Davies et al. (2004) Davies, J., Minchin, R., Sabatini, S., et al. 2004, MNRAS, 349, 922
- Davies et al. (2006) Davies, J. I., Disney, M. J., Minchin, R. F., Auld, R., & Smith, R. 2006, MNRAS, 368, 1479
- Dey et al. (2019) Dey, A., Schlegel, D. J., Lang, D., et al. 2019, AJ, 157, 168
- Dey et al. (2025) Dey, S., Jones, M. G., Sand, D. J., et al. 2025, ApJ, 983, 2
- Di Teodoro & Fraternali (2015) Di Teodoro, E. M. & Fraternali, F. 2015, MNRAS, 451, 3021
- Duc & Bournaud (2008) Duc, P.-A. & Bournaud, F. 2008, ApJ, 673, 787
- Efstathiou (1992) Efstathiou, G. 1992, MNRAS, 256, 43P
- Ferrarese et al. (2012) Ferrarese, L., Côté, P., Cuillandre, J.-C., et al. 2012, ApJS, 200, 4
- Gavazzi et al. (2005) Gavazzi, G., Boselli, A., van Driel, W., & O’Neil, K. 2005, A&A, 429, 439
- Giovanelli et al. (2005) Giovanelli, R., Haynes, M. P., Kent, B. R., et al. 2005, AJ, 130, 2598
- Guo et al. (2021) Guo, H., Jones, M. G., Wang, J., & Lin, L. 2021, ApJ, 918, 53
- Haynes et al. (2007) Haynes, M. P., Giovanelli, R., & Kent, B. R. 2007, ApJ, 665, L19
- Haynes et al. (2018) Haynes, M. P., Giovanelli, R., Kent, B. R., et al. 2018, ApJ, 861, 49
- Heyer et al. (2009) Heyer, M., Krawczyk, C., Duval, J., & Jackson, J. M. 2009, ApJ, 699, 1092
- Hoeft & Gottlöber (2010) Hoeft, M. & Gottlöber, S. 2010, Advances in Astronomy, 2010, 693968
- Hoeft et al. (2006) Hoeft, M., Yepes, G., Gottlöber, S., & Springel, V. 2006, MNRAS, 371, 401
- Hoffman et al. (1987) Hoffman, G. L., Helou, G., Salpeter, E. E., Glosson, J., & Sandage, A. 1987, ApJS, 63, 247
- Hoffman et al. (1996) Hoffman, G. L., Salpeter, E. E., Farhat, B., et al. 1996, ApJS, 105, 269
- Huang et al. (2012) Huang, S., Haynes, M. P., Giovanelli, R., & Brinchmann, J. 2012, ApJ, 756, 113
- Hunter et al. (2012) Hunter, D. A., Ficut-Vicas, D., Ashley, T., et al. 2012, AJ, 144, 134
- Janowiecki et al. (2015) Janowiecki, S., Leisman, L., Józsa, G., et al. 2015, ApJ, 801, 96
- Jiang et al. (2019) Jiang, P., Yue, Y., Gan, H., et al. 2019, Science China Physics, Mechanics, and Astronomy, 62, 959502
- Jing et al. (2024) Jing, Y., Wang, J., Xu, C., et al. 2024, Science China Physics, Mechanics, and Astronomy, 67, 259514
- Jones et al. (2022a) Jones, M. G., Sand, D. J., Bellazzini, M., et al. 2022a, ApJ, 926, L15
- Jones et al. (2022b) Jones, M. G., Sand, D. J., Bellazzini, M., et al. 2022b, ApJ, 935, 51
- Junais et al. (2021) Junais, Boissier, S., Boselli, A., et al. 2021, A&A, 650, A99
- Junais et al. (2022) Junais, Boissier, S., Boselli, A., et al. 2022, A&A, 667, A76
- Kashibadze et al. (2020) Kashibadze, O. G., Karachentsev, I. D., & Karachentseva, V. E. 2020, A&A, 635, A135
- Kennicutt (1998) Kennicutt, Robert C., J. 1998, ApJ, 498, 541
- Kennicutt et al. (2011) Kennicutt, R. C., Calzetti, D., Aniano, G., et al. 2011, PASP, 123, 1347
- Kent et al. (2008) Kent, B. R., Giovanelli, R., Haynes, M. P., et al. 2008, AJ, 136, 713
- Klypin et al. (1999) Klypin, A., Kravtsov, A. V., Valenzuela, O., & Prada, F. 1999, ApJ, 522, 82
- Koribalski et al. (2020) Koribalski, B. S., Staveley-Smith, L., Westmeier, T., et al. 2020, Ap&SS, 365, 118
- Larson (1981) Larson, R. B. 1981, MNRAS, 194, 809
- Lee et al. (2024) Lee, G., Hwang, H. S., Lee, J., Shin, J., & Song, H. 2024, ApJ, 962, 129
- Lee et al. (2022) Lee, J., Kimm, T., Blaizot, J., et al. 2022, ApJ, 928, 144
- Leisman et al. (2021) Leisman, L., Rhode, K. L., Ball, C., et al. 2021, AJ, 162, 274
- Lelli (2022) Lelli, F. 2022, Nature Astronomy, 6, 35
- Longobardi et al. (2020) Longobardi, A., Boselli, A., Fossati, M., et al. 2020, A&A, 644, A161
- Mamon et al. (2010) Mamon, G. A., Biviano, A., & Murante, G. 2010, A&A, 520, A30
- McMullin et al. (2007) McMullin, J. P., Waters, B., Schiebel, D., Young, W., & Golap, K. 2007, in Astronomical Society of the Pacific Conference Series, Vol. 376, Astronomical Data Analysis Software and Systems XVI, ed. R. A. Shaw, F. Hill, & D. J. Bell, 127
- Minchin et al. (2005) Minchin, R., Davies, J., Disney, M., et al. 2005, ApJ, 622, L21
- Minchin et al. (2007) Minchin, R., Davies, J., Disney, M., et al. 2007, ApJ, 670, 1056
- Montes et al. (2024) Montes, M., Trujillo, I., Karunakaran, A., et al. 2024, A&A, 681, A15
- Nan et al. (2011) Nan, R., Li, D., Jin, C., et al. 2011, International Journal of Modern Physics D, 20, 989
- O’Beirne et al. (2024) O’Beirne, T., Staveley-Smith, L., Wong, O. I., et al. 2024, MNRAS, 528, 4010
- Papastergis et al. (2016) Papastergis, E., Adams, E. A. K., & van der Hulst, J. M. 2016, A&A, 593, A39
- Ponomareva et al. (2021) Ponomareva, A. A., Mulaudzi, W., Maddox, N., et al. 2021, MNRAS, 508, 1195
- Qian et al. (2020) Qian, L., Yao, R., Sun, J., et al. 2020, The Innovation, 1, 100053
- Roediger & Brüggen (2007) Roediger, E. & Brüggen, M. 2007, MNRAS, 380, 1399
- Sardaneta et al. (2022) Sardaneta, M. M., Amram, P., Boselli, A., et al. 2022, A&A, 659, A45
- Schiminovich et al. (2010) Schiminovich, D., Catinella, B., Kauffmann, G., et al. 2010, MNRAS, 408, 919
- Serra et al. (2024) Serra, P., Oosterloo, T. A., Kamphuis, P., et al. 2024, A&A, 690, A4
- Smith et al. (2015) Smith, R., Flynn, C., Candlish, G. N., Fellhauer, M., & Gibson, B. K. 2015, MNRAS, 448, 2934
- Taylor & Webster (2005) Taylor, E. N. & Webster, R. L. 2005, ApJ, 634, 1067
- Thoul & Weinberg (1996) Thoul, A. A. & Weinberg, D. H. 1996, ApJ, 465, 608
- Tonnesen & Bryan (2021) Tonnesen, S. & Bryan, G. L. 2021, ApJ, 911, 68
- Wang et al. (2016) Wang, J., Koribalski, B. S., Serra, P., et al. 2016, MNRAS, 460, 2143
- Wang et al. (2023) Wang, J., Yang, D., Oh, S. H., et al. 2023, ApJ, 944, 102
- Wong & Blitz (2002) Wong, T. & Blitz, L. 2002, ApJ, 569, 157
- Xu et al. (2025) Xu, C., Wang, J., Jing, Y., et al. 2025, Research in Astronomy and Astrophysics, 25, 015011
- Xu et al. (2023) Xu, J.-L., Zhu, M., Yu, N., et al. 2023, ApJ, 944, L40
- Yang et al. (2025) Yang, D., Wang, J., Qu, Z., et al. 2025, ApJ, 984, 15
- Yu et al. (2024) Yu, N.-P., Zhu, M., Xu, J.-L., et al. 2024, MNRAS, 532, 1744
- Zaritsky et al. (2023) Zaritsky, D., Donnerstein, R., Dey, A., et al. 2023, ApJS, 267, 27
- Zhang et al. (2024) Zhang, C.-P., Zhu, M., Jiang, P., et al. 2024, Science China Physics, Mechanics, and Astronomy, 67, 219511
- Zhu et al. (2021) Zhu, M., Yu, H., Wang, J., et al. 2021, ApJ, 922, L21
Appendix A PV diagram of VLA observation for AGC 226178